# Large Language Models Can Help Mitigate Barren Plateaus

### v1

@@ 3 Experiment
We analyze the sensitivity of hyperparameters, including Temperature and Top P, for llms. Temperature controls the randomness of predictions, with higher values generating more diverse outputs, while Top P affects the probabilities of token selections, ensuring a more focused yet flexible generation. To identify the optimal settings, we first narrowed down the hyperparameter ranges through manual tuning and then applied grid search to determine the best combinations (Temperature, Top P) for each dataset: Iris (0.5, 0.9), Wine (0.1, 0.45), Titanic (0.8, 0.75), and MNIST (0.8, 0.8), as presented in Fig. 7. The combinations of the above hyperparameters were subsequently used in this study.
@@
Analysis of the expected improvement.
We analyze the patterns on the expected improvement (EI) and the corresponding gradient variance across various qnn structures as search iterations progress. Representative experiments conducted on Iris are illustrated in Fig. 8 as an example. Our findings show that as the model size grows, more search iterations are required to obtain optimal initial parameters that enable qnns to maintain higher gradient variance in training. This is expected, as larger models expand the search space, demanding greater computational resources to explore effectively.
4 Related Work
McClean et al. (2018) first investigated barren plateau (BP) phenomenons and demonstrated that under the assumption of the 2-design Haar distribution, gradient variance in qnns will exponentially decrease to zero during training as the model size increases. In recent years, enormous studies have been devoted to mitigating BP issues in qnnQi et al. (2023). Cunningham and Zhuang (2024) categorize most existing studies into the following five groups. (i) Initialization-based strategies initialize model parameters with various well-designed distributions in the initialization stage Grant et al. (2019); Sack et al. (2022); Mele et al. (2022); Grimsley et al. (2023); Liu et al. (2023); Park and Killoran (2024). (ii) Optimization-based strategies address BP issues and further enhance trainability during optimization Ostaszewski et al. (2021); Suzuki et al. (2021); Heyraud et al. (2023); Liu et al. (2024); Sannia et al. (2024). (iii) Model-based strategies attempt to mitigate BPs by proposing new model architectures Li et al. (2021); Bharti and Haug (2021); Du et al. (2022); Selvarajan et al. (2023); Tüysüz et al. (2023); Kashif and Al-Kuwari (2024). (iv) To address both BPs and saddle points, Zhuang et al. (2024) regularize qnns’ model parameters via Bayesian approaches. (v) Rappaport et al. (2023) measure BP phenomenon via various informative metrics.
5 Conclusion
In this study, we proposed a new llm-driven framework, AdaInit, designed to mitigate barren plateaus (BPs) in qnn’s training. By iteratively refining qnn’s initialization through adaptive prompting and posterior estimation, AdaInit can maximize gradient variance, improving qnn’s trainability against BPs. Our theoretical analysis establishes the submartingale property of expected improvement (EI), ensuring the iterative search can eventually identify optimal initial model parameters for qnn. Through extensive experiments across four public datasets, we demonstrated that AdaInit outperforms conventional classic initialization methods in maintaining higher gradient variance as qnn’s sizes increase. Overall, this study paves a new way to explore how llms help mitigate BPs in qnn’s training.
Limitations & future work.
First, in our theoretical analysis, we assume that the maximum gradient of qnns is bounded by a positive constant, i.e., the gradient doesn’t explode during training. This assumption is practical in most cases. Besides, we rigorously prove that the submartingale has a supremum in our settings. In the future, we plan to prove that convergence of submartingale is guaranteed in a finite number of search iterations.
Appendix A APPENDIX
In the appendix, we present the architecture of the quantum circuit and our hardware/software. Besides, we display the prompt designs in this study.
Model architecture of the quantum circuit.
In this study, we examine our proposed framework using a backbone qnn, which concatenates the following quantum circuit with a fully connected layer. The circuit architecture is described in Figure 9.
Hardware and software.
The experiment is conducted on a server with the following settings:
  • Operating System: Ubuntu 22.04.3 LTS
  • CPU: Intel Xeon w5-3433 @ 4.20 GHz
  • GPU: NVIDIA RTX A6000 48GB
  • Software: Python 3.11, PyTorch 2.1, Pennylane 0.31.1.
@@ Appendix A APPENDIX
Before presenting the prompts, we first introduce the notation for the hyperparameter in prompts. ‘nlayers’, ‘nqubits’, ‘nrot’, ‘nclasses’ denote the number of layers, qubits, rotation gates, and classes for the qnn, respectively. ‘init’ denotes the initial data distribution for the qnn. ‘data_desc’ denotes the data description. ‘feedback’ denotes the gradient feedback from the previous iteration during the search.
Proof of Lemmas.
We provide a rigorous proof of the following lemmas.
Lem. 1.
then the gap between $\partial E^{(t)}$ and $\overline{\partial E}$ will not exceed the range of $[\partial E_{min},\partial E_{max}]$:
Thus, we have:
$\leq\frac{1}{T_{tr}}\sum_{t=1}^{T_{tr}}(\partial E_{max}-\partial E_{min})^{2}$
$=(\partial E_{max}-\partial E_{min})^{2}.$
Lem. 2.
From Def. 1, for $\forall$ $t\in\mathbb{Z}^{+}$, in the $t$-th search iteration, we have:
$\Delta^{(t)}=\max(\mathrm{Var}[\partial E^{(t)}]-S^{(t-1)},0).$
Combining with Lem. 1, for $\forall$ $t\in\mathbb{Z}^{+}$, we have:
which indicates that:
$\Delta^{(t)}\leq(\partial E_{max}-\partial E_{min})^{2}.$
Lem. 3.
A process $S^{(t)}$ is submartingale relative to $(\Omega,\mathcal{F},P)$ if the following three conditions, Adaptedness, Integrability, and Submartingale condition, hold true Williams (1991).
Adaptedness.
We first aim to verify that $S^{(t)}$ is determined based on the information available up to past $t$ iterations. By Def. 1, $S^{(t)}=\sum_{t_{i}=1}^{t}\Delta^{(t_{i})}\cdot I^{(t_{i})}$ is a finite sum of random variables that are measurable w.r.t. $\sigma\bigl{(}\Delta^{(1)},\Delta^{(2)},\dots,\Delta^{(t)}\bigr{)}$. Thus, $S^{(t)}$ is also measurable w.r.t. $\mathcal{F}^{(t)}$, ensuring the adaptedness.
Integrability.
In Lem. 2, $\Delta^{(t)}\leq(\partial E_{max}-\partial E_{min})^{2}$ for $\forall$ $t\in\mathbb{Z}^{+}$. Thus,
<<$\infty$,< ∞ ,
which ensures $\mathbb{E}[|S^{(t)}|]$ is integrable for each $t$.
Submartingale condition.
We observe that
$S^{(t)}=S^{(t-1)}+\Delta^{(t)}.$
Thus, given $\mathcal{F}^{(t-1)}$, we have
$\mathbb{E}\bigl{[}S^{(t)}\big{|}\mathcal{F}^{(t-1)}\bigr{]}=\mathbb{E}\bigl{[}S^{(t-1)}+\Delta^{(t)}\big{|}\mathcal{F}^{(t-1)}\bigr{]},$
$\mathbb{E}\bigl{[}S^{(t)}\big{|}\mathcal{F}^{(t-1)}\bigr{]}$ $=\mathbb{E}\bigl{[}S^{(t-1)}+\Delta^{(t)}\big{|}\mathcal{F}^{(t-1)}\bigr{]}$
$=S^{(t-1)}+\mathbb{E}[\Delta^{(t)}]$
$=S^{(t-1)}+\bigl{(}\delta p+0(1-p)\bigr{)}$
$\geq S^{(t-1)},$
where $\delta$ denotes a positive increment when $\Delta^{(t)}>\frac{1}{poly(N,L)T}$.
Thus, the submartingale condition holds true for $\forall$ $t\geq 1$ s.t.
$\mathbb{E}\bigl{[}S^{(t)}\big{|}\mathcal{F}^{(t-1)}\bigr{]}\geq S^{(t-1)},\quad\forall t\geq 1,$
Lem. 4.
Since the process $\{S^{(t)}\}_{t\geq 1}$ is a $L^{1}$-bounded submartingale s.t. $\sup_{t}\mathbb{E}[|S^{(t)}|]<\infty$, we apply Doob’s Forward Convergence Theorem Williams (1991), which guarantees the almost sure existence of a finite random variable $S^{(\infty)}$ s.t. $S^{(\infty)}$=$\lim_{t\to\infty}S^{(t)}$. This implies that the process $\{S^{(t)}\}$ has a well-defined almost sure limit.
Computational budgets.
Based on the above computing infrastructure and settings, computational budgets in our experiments are described as follows. Our search framework can be reproduced within one hour given that the number of qubits is less than 18. The total experiments take around 600 hours to complete.
Ethical and broader impacts.
We confirm that we fulfill the author’s responsibilities and address the potential ethical issues. This work paves a novel way to explore how generative models, such as llms, help improve the trainability of qnns, which could benefit the community of natural language processing and quantum machine learning.
Statement of data privacy.
The datasets used in this study were obtained from publicly available sources.
Potential risk.
Reproducing the experiments in this study may take significant time and computational resources.
Disclaimer regarding human subjects results.
NA

### v2

@@ 5 Experiment
We analyze the sensitivity of hyperparameters, including Temperature and Top P, for llms. Temperature controls the randomness of predictions, with higher values generating more diverse outputs, while Top P affects the probabilities of token selections, balancing generation diversity and structural consistency. To identify optimal settings, we first narrowed down the ranges through manual tuning and then applied grid search to determine the best combinations (Temperature, Top P) for each dataset: Iris (0.5, 0.9), Wine (0.1, 0.45), Titanic (0.8, 0.75), and MNIST (0.8, 0.8), as presented in Fig. 7. The combinations of the above hyperparameters were used in this study.
Due to the page limit, we present our supplementary results, such as simulation time of qnns, and the computational trade-off, in the Appendix B.
6 Conclusion
In this study, we aim to mitigate barren plateaus (BPs) by introducing a new AI-driven submartingale-based framework, namely AdaInit. This framework iteratively generates effective initial parameters using generative models, such as llms, for qnns that yield non-negligible gradient variance, thereby mitigating BPs. Our theoretical analysis establishes the submartingale property for the iterative process, ensuring the effective generation. Through extensive experiments across various model scales, we demonstrated that AdaInit outperforms conventional classic initialization methods in maintaining higher gradient variance as qnn’s sizes increase. Overall, this study might initiate a new avenue to explore how llms help mitigate BPs.
Limitations, future work, and broad impact.
First, our theoretical analyses assume that the maximum gradient of qnns is bounded by a positive constant, implying that gradient explosion does not occur during training, a condition that is typically satisfied in practice. Second, due to the practical limitations of quantum simulation, our experiments are constrained to qnns with up to 20 qubits. We also assume an idealized setting where quantum measurements are noise-free. Moreover, our current scope excludes ansatz-induced BPs, which may be mitigated through architectural modifications, as discussed above. For future work, we plan to (i) accelerate the convergence of the iterative process and (ii) expand the applicability of our framework beyond BP mitigation. In particular, it can be leveraged to guide qnn architecture design or to identify optimal model parameters in training. More broadly, our framework can support robust qnn training across various domains, such as healthcare, where robustness and reliability are critical.
Ethics Statement
This work does not involve human subjects, sensitive personal data, or tasks with foreseeable negative societal impact. The datasets used (Iris, Wine, Titanic, and MNIST) are standard, publicly available benchmarks. We ensured compliance with their respective licenses and data usage guidelines. The proposed framework is designed to mitigate barren plateaus and does not directly enable harmful applications. Nonetheless, as with other advances in optimization, the method could potentially be applied in sensitive domains; in such cases, practitioners should carefully consider fairness, privacy, and security concerns in line with the ICLR Code of Ethics.
Reproducibility Statement
We have made every effort to ensure reproducibility of our results. Detailed theoretical proofs are included in Appendix A, whereas experimental settings, dataset splits, hyperparameters, architecture of the backbone quantum circuit, and the computing infrastructure are described in Appendix B. Besides, we provide sample code and datasets in the supplementary materials, which will be made publicly available upon publication. These resources, together with the descriptions of prompts, model parameters, and evaluation protocols, are intended to enable independent verification and extension of our work.
Appendix A Proofs
In this section, we provide formal proofs that consolidate the theoretical guarantees of our framework.
Proof for Lemma 2.
then the gap between $\partial E^{(t)}$ and $\overline{\partial E}$ will not exceed the range of $\partial E_{min},\partial E_{max}]$:
Thus, we have:
$\leq\frac{1}{T_{tr}}\sum_{t=1}^{T_{tr}}(\partial E_{max}-\partial E_{min})^{2}$
$=(\partial E_{max}-\partial E_{min})^{2}.$
Proof for Lemma 3.
From Def. 2, for $\forall$ $t\in\mathbb{Z}^{+}$, in the $t$-th search iteration, we have:
$\Delta^{(t)}=\max(\mathrm{Var}[\partial E^{(t)}]-S^{(t-1)},0).$
Combining with Lem. 2, for $\forall$ $t\in\mathbb{Z}^{+}$, we have:
which indicates that:
$\Delta^{(t)}\leq(\partial E_{max}-\partial E_{min})^{2}.$
Before introducing the formal statement of submartingale property, we first define the random variables. Let $\alpha=\nicefrac{{1}}{{(poly(N,L)K)}}$. To avoid confusion, notably, we clarify that the $K$ refers to a previously defined parameter, whereas in Tab. 1, the suffix ‘K’ attached to numbers represents the unit ‘thousand’. Formally, we define discrete random variables $I^{(t)}$ and continuous random variables $\Delta^{{(t)}}$ as follows:
$\begin{cases}P(I^{(t)}=1)=P(\Delta^{(t)}\geq\alpha)=p,\\ P(I^{(t)}=0)=P(\Delta^{(t)}<\alpha)=1-p,\end{cases}$ (5)
with a real number $p\in(0,1]$. Hence, $I^{(t)}$ is a Bernoulli random variable as an indicator of the event $\Delta^{(t)}\geq\alpha$, and its associated continuous random variable $\Delta^{{(t)}}$ is implicitly defined with an arbitrary probability density function $p(y)$ that satisfies
$P(\Delta^{(t)}\geq\alpha)=\int_{\alpha}^{\infty}p(y)dy=p.$ (6)
With such a relationship between them, the following can be easily verified, for all $t\in\mathbb{Z}^{+}$,
$P(I^{(t)}=x|\Delta^{(t)}=y)=\begin{cases}x,&y\geq\alpha\\ 1-x,&\hbox{o.w.}\end{cases}.$ (7)
Lemma 6 (Submartingale Property, formal statement of Lemma 4).
Let $\{I^{(t)},\Delta^{(t)}\}_{t\geq 1}$ be a sequence of i.i.d. random variables as defined in Eq. 5. We define joint random variables $W^{(t)}=\Delta^{(t)}\cdot I^{(t)}$ and the natural filtration $\mathcal{F}^{(t)}=\sigma\bigl(W^{(1)},\cdots,W^{(t)}\bigr)$. Then the sequence $\{S^{(t)}\}_{t\geq 1}$ as defined by Def. 2 is a submartingale with respect to the filtration $\{\mathcal{F}^{(t)}\}_{t\geq 1}$.
Proof for Lemma 6.
According to Def. 1, a process $S^{(t)}$ is a submartingale relative to $(\Omega,\mathcal{F},P)$ if it satisfies Adaptedness, Integrability, and Submartingale.
Adaptedness.
We first aim to verify that $S^{(t)}$ is determined based on the information available up to past $t$ iterations. By Def. 2, $S^{(t)}=\sum_{t_{i}=1}^{t}\Delta^{(t_{i})}\cdot I^{(t_{i})}=\sum_{t_{i}=1}^{t}W^{(t_{i})}$ is a finite sum of random variables that are measurable w.r.t. $\sigma\bigl(W^{(1)},\dots,W^{(t)}\bigr)$ (or $\sigma\bigl(I^{(1)},\dots,I^{(t)}\bigr)$ for short due to the relationship between $I^{(t)}$ and $\Delta^{(t)}$ as shown in Eq. 5). Thus, $S^{(t)}$ is also measurable w.r.t. $\mathcal{F}^{(t)}$, ensuring the adaptedness.
Integrability.
In Lem. 3, $\Delta^{(t)}\leq(\partial E_{max}-\partial E_{min})^{2}$ for $\forall$ $t\in\mathbb{Z}^{+}$. Thus,
$<\infty,$
which ensures $\mathbb{E}[|S^{(t)}|]$ is integrable for each $t$.
Submartingale.
Before proving this condition, we show the following necessary inequality: with $\alpha=\nicefrac{{1}}{{(poly(N,L)K)}}$,
$\mathbb{E}[\Delta^{(t)}\cdot I^{(t)}]$ $=\sum_{x=0,1}\int_{-\infty}^{\infty}P(I^{(t)}=x|\Delta^{(t)}=y)\cdot p(y)\cdot x\cdot y\ dy$
$=\int_{\alpha}^{\infty}p(y)\cdot y\ \ dy$
$\geq\alpha\int_{\alpha}^{\infty}p(y)dy$
$=\alpha p\ ,$ (8)
where the second step follows from Eq. 7 and the last step uses Eq. 6. Apparently $\alpha p>0$.
We observe that
$S^{(t)}=S^{(t-1)}+\Delta^{(t)}\cdot I^{(t)}.$
$\mathbb{E}\bigl[S^{(t)}\big|\mathcal{F}^{(t-1)}\bigr]$ $=\mathbb{E}\bigl[S^{(t-1)}+\Delta^{(t)}\cdot I^{(t)}\big|\mathcal{F}^{(t-1)}\bigr]$
$=S^{(t-1)}+\mathbb{E}[\Delta^{(t)}\cdot I^{(t)}]$
$\geq S^{(t-1)}+\alpha p$
$\geq S^{(t-1)},$
where the last two step applies Appendix A.
Thus, the submartingale condition holds true for $\forall$ $t\geq 1$ s.t.
$\mathbb{E}\bigl[S^{(t)}\big|\mathcal{F}^{(t-1)}\bigr]\geq S^{(t-1)},\quad\forall\ t\geq 1.$
∎ Intuitively, $S^{(t)}$ tracks the cumulative amount of non-trivial variance improvements observed over the iterations — akin to measuring meaningful progress in exploration.
Proof for Lemma 5.
Since the process $\{S^{(t)}\}_{t\geq 1}$ is a $L^{1}$-bounded submartingale s.t. $\sup_{t}\mathbb{E}[|S^{(t)}|]<\infty$, we apply the Doob’s Forward Convergence Theorem (by Thm. 1), which guarantees the almost sure existence of a finite random variable $S^{(\infty)}$ s.t. $S^{(\infty)}=\lim_{t\to\infty}S^{(t)}$. This implies that the process $\{S^{(t)}\}$ has a well-defined almost sure limit.
Proof for Theorem 4.
To analyze the expected hitting time, we first construct a drift-adjusted process $\{Z^{(t)}\}_{t\geq 1}$ adapted to a filtration $\{\mathcal{F}^{(t)}\}_{t\geq 1}$ as $Z^{(t)}=S^{(t)}-\delta t$, where $\delta=\nicefrac{{p}}{{(poly(N,L)K)}}>0$.
Given $\alpha=\nicefrac{{1}}{{(poly(N,L)K)}}$ and follow those similar steps in the proof for Lem. 6, we can derive
$\mathbb{E}[\Delta^{(t)}\cdot I^{(t)}]\geq\alpha p=\delta,$ (9)
where the last step is by definition of $\delta$.
We then verify that $Z^{(t)}$} is also a submartingale:
Adaptedness: Similr to $S^{(t)}$, $Z^{(t)}$} is also determined by the past $t$ iterations w.r.t. the same filtration $\sigma\bigl(W^{(1)},\dots,W^{(t)}\bigr)$ as $S^{(t)}$. Thus, $Z^{(t)}$} can meet the adaptedness.
Integrability: Given that $S^{(t)}$ is $L^{1}$-bounded, and $Z^{(t)}$ is obtained by subtracting a deterministic finite value $\delta t$ from $S^{(t)}$, it follows immediately that $Z^{(t)}$ is also $L^{1}$-bounded, i.e., $\mathbb{E}[|Z^{(t)}|]$ is integrable for each $t$.
Submartingale: We further show that $Z^{(t)}$ meets the submartingale inequality as follows.
$\mathbb{E}[Z^{(t+1)}\mid\mathcal{F}^{(t)}]$ $=\mathbb{E}[S^{(t+1)}-\delta(t+1)\mid\mathcal{F}^{(t)}]$
$=\mathbb{E}[S^{(t+1)}\mid\mathcal{F}^{(t)}]-\delta(t+1)$
$=\mathbb{E}[S^{(t)}+\Delta^{(t+1)}\cdot I^{(t+1)}\mid\mathcal{F}^{(t)}]-\delta(t+1)$
$=S^{(t)}+\mathbb{E}[\Delta^{(t+1)}\cdot I^{(t+1)}\mid\mathcal{F}^{(t)}]-\delta(t+1)$
$\geq S^{(t)}+\delta-\delta(t+1)$
$=Z^{(t)},$
where the fifth step is obtained by combining the fact that $\Delta^{(t+1)}\cdot I^{(t+1)}$ is independent of $\mathcal{F}^{(t)}$ and Eq. 9.
Second, we define the hitting time $T_{b}$ as $T_{b}=\inf\left\{T\in\mathbb{Z}^{+}:S^{(T)}=b\right\}$. Without loss of generality, we assume $b\leq B_{S}$ since $S^{(t)}\leq B_{S}$ a.s., for $\forall\ t\in\mathbb{Z}^{+}$ (by Lem. 5). We further verify that $T_{b}$ is a bounded stopping time as follows.
We observe that $\{T_{b}\leq T\}=\{\exists\ t\leq T\hbox{ such that }S^{(t)}=b\}=\bigcup_{t=0}^{T}\{S^{(t)}=b\}$. Since $S^{(t)}$ is $\mathcal{F}^{(T)}$ measurable for $\forall\ t\leq T$, we have $\{S^{(t)}=b\}\in\mathcal{F}^{(T)}$, which indicates that the finite union $\bigcup_{t=0}^{T}\{S^{(t)}=b\}\in\mathcal{F}^{(T)}$. Hence, $\{T_{b}\leq T\}\in\mathcal{F}^{(T)}$, which by definition shows that $T_{b}$ is a bounded stopping time.
Third, we define $T_{b}\wedge t$ as $\min(T_{b},\ t)$. Given that $T_{b}$ is a bounded stopping time, $T_{b}\wedge t$ is also a bounded stopping time (by Lem. 1). Based on this condition, the Doob’s Optional Stopping Theorem implies that $\mathbb{E}[Z^{(T_{b}\wedge t)}]\geq\mathbb{E}[Z^{(0)}]=0$ (by Thm. 2). Thus, we have:
$\mathbb{E}[S^{(T_{b}\wedge t)}-\delta(T_{b}\wedge t)]\geq 0$
implying
$\mathbb{E}[S^{(T_{b}\wedge t)}]\geq\delta\mathbb{E}[T_{b}\wedge t].$
Since $S^{(t)}$ is non-decreasing and bounded by $B_{S}$, we have $T_{b}\wedge t=T_{b}$ as $t\to\infty$ almost surely, which implies that $\mathbb{E}[T_{b}\wedge t]=\mathbb{E}[T_{b}]$. Moreover, by the definition of $T_{b}$, it follows that $S^{(T_{b}\wedge t)}\to b$ as $t\to\infty$, i.e., $\mathbb{E}[S^{(T_{b}\wedge t)}]$ is bounded by a dominating constant $b$. So, by the Dominated Convergence Theorem, as $t\to\infty$, we have $\mathbb{E}[S^{(T_{b}\wedge t)}]\to\mathbb{E}[S^{(T_{b})}]=b$ (by Thm. 3).
By integrating the above equations and taking the limit, we conclude:
$\mathbb{E}[T_{b}]\leq\frac{b}{\delta}=\frac{bK\cdot poly(N,L)}{p}.$
Thus, the following result can be derived immediately by plugging concrete values of $b$ into Thm. 4.
Corollary 1 (Expected Hitting Time Under Specific Thresholds).
  1. 1.
    With an expected number of $\nicefrac{{K}}{{p}}$ iterations, Algo. 1 can identify a candidate model parameter $\bm{\theta}^{*}_{0}$ that has $\mathrm{Var}[\partial E]\approx\nicefrac{{1}}{{poly(N,L)}}$.
  2. 2.
    With an expected number of $\nicefrac{{B_{S}\cdot K\cdot poly(N,L)}}{{p}}$ iterations, Algo. 1 can identify a candidate model parameter $\bm{\theta}^{*}_{0}$ that has $\mathrm{Var}[\partial E]=\mathcal{O}(poly(N,L))$.
Appendix B Supplementary Experiments
In this section, we present supplementary details about our experimental results.
Dataset $\left|D\right|$ $\left|F\right|$ $\left|C\right|$ Splits
Iris 150 4 3 60:20:20
Wine 178 13 3 80:20:30
Titanic 891 11 2 320:80:179
MNIST 60,000 784 10 320:80:400
Dataset.
We evaluate our proposed method across four public datasets that are widely used in quantum machine learning. Iris  [^1] https://archive.ics.uci.edu/ml/datasets/iris (Fisher, 1936) is a classic machine-learning benchmark that measures various attributes of three-species iris flowers. Wine  [^2] https://archive.ics.uci.edu/ml/datasets/iris (Fisher, 1936) is a well-known dataset that includes 13 attributes of chemical composition in wines. Titanic  [^3] https://www.kaggle.com/c/titanic (Kaggle, 2012) contains historical data about passengers aboard the Titanic and is typically used to predict survival. MNIST  [^4] http://yann.lecun.com/exdb/mnist/ (LeCun et al., 2010) is a widely used small benchmark in computer vision. This benchmark consists of $\times28$ gray-scale images of handwritten digits from 0 to 9. We follow the settings of BeInit Kulshrestha & Safro (2022) and conduct experiments in binary classification. Specifically, we randomly sub-sample a certain number of instances from the first two classes of each dataset to create a new subset. After sub-sampling, we employ the t-SNE technique to reduce the feature dimensions to ensure they do not exceed the number of available qubits. The statistics of the original datasets, along with the data splits for training, validation, and testing, are presented in Table 2. Importantly, the total number of sub-sampled instances corresponds to the sum of the split datasets. For instance, in the Iris dataset, the total number of sub-sampled instances is 100.
Figure 8: Comparison of model parameter initializations using a classic method, random initializer (RI), and llms. All methods apply a uniform distribution.
Contribution of llms.
Besides comparing our framework with the classic method, we further investigate llms’ contribution to the initialization process on Iris. Specifically, we compare the generator within our framework when initialized using a random initializer (RI), which uniformly generates the parameters, versus using llm-based uniform initialization (llms). As shown in Fig. 8, “RI” performs comparably to, or even worse than, the “Classic” method (uniform), as it generates random parameters without any guided refinement. In contrast, “llms” consistently outperforms both baselines, achieving significantly higher gradient variance. These findings suggest that the llm-driven generator can more effectively explore the parameter space and identify better initializations within a limited number of iterations.
LLaMA 3 Variables Layer Expected Actual
70B $N\in[2,20]$ 0 (2, $N$, 3) (2, 3)
70B $L\in[4,40]$ 0 ($L$, 2, 3) ($L$, 3)
405B $L\in[4,40]$ 1 (2, 2) (2, 2, 2)
Table 3: Comparison of generated model parameters (layer, expected shape, and actual shape) between two open-source llms—LLaMA 3 70B Instruct and LLaMA 3 405B Instruct—evaluated under various numbers of qubits (NN) or layers ($L$).
Observations of open-source models.
We observe that two LLaMA 3 open-source llmGrattafiori et al. (2024) perform significantly worse. To further understand these failures, we present representative cases in Tab. 3 and observe that both models fail to generate correct shapes of model parameters, indicating that these open-source models struggle to follow precise structural instructions and suggesting their current limitations in shape-constrained generation tasks.
Figure 9: Assessment of the simulation time in qnn training.
Simulation time in qnn training.
We assess the simulation time of qnn training under varying model sizes (number of qubits, $N\in[2,20]$) and subsampled MNIST dataset sizes (number of instances, $|D|\in[800,4000]$). We train QNNs for 30 epochs and present the average runtime per epoch. When varying $N$, we fix the number of layers $L$ at 2; when varying $|D|$, we fix both $N$ and $L$ as 2. As presented in Fig. 9, in a classical simulated environment, the average training time of qnns increases exponentially w.r.t. the number of qubits, while it grows roughly linearly with the dataset size. These observations reflect an inherent scalability issue in classical simulation of quantum systems. Such limitations are widely acknowledged in the quantum computing community and are unlikely to be fully overcome until practical quantum hardware becomes more accessible.
Figure 10: Trade-off analysis between computational cost and performance benefits for 2 qubits (left) and 20 qubits (right) setups.
Trade-off analysis.
To analyze the trade-off in Fig. 11, we present the relationship between computational cost (measured by the number of search iterations) and performance benefits (quantified by gradient variance) in Fig. 10. We observe a roughly linear relationship between them. In the 2-qubit case, 40% cost yields over 60% gain, while in the 20-qubit case, 35% cost yields over 44% gain, showing strong early-stage cost-effectiveness. The diminishing returns after a few iterations align with submartingale optimization behavior, and the consistent trends across scales highlight AdaInit’s suitability for budget-aware scenarios.
@@ Appendix B Supplementary Experiments
Patterns of Expected Improvement (EI).
Due to limited pages, we present the patterns of EI in Fig. 11.
Figure 12: Trade-off analysis of the assumed lower bound, comparing polynomial terms against the exponential baseline, shown on a log scale.
Empirical analysis of the assumed lower bound.
To determine the assumed lower bound, $\nicefrac{{1}}{{(poly(N,L)K)}}$, we conduct a trade-off analysis. A larger polynomial coefficient enlarges the admissible regime, but at the cost of including cases with vanishingly small gradient variance, whereas a smaller coefficient may filter out meaningful expected improvements, thereby preventing the framework from effectively exploring initial parameters. To proactively mitigate BPs, we restrict our attention to the range of qubits that are particularly susceptible to BPs. As illustrated in Fig. 12, we vary the polynomial coefficient and compare against the exponential baseline $\nicefrac{{1}}{{(K2^{2N})}}$. Considering the trade-off, we empirically select $\nicefrac{{1}}{{(KN^{6})}}$ as the lower bound.
@@
Before presenting the prompts, we first introduce the notation for the hyperparameter in the prompts. ‘nlayers’, ‘nqubits’, ‘nrot’, ‘nclasses’ denote the number of layers, qubits, rotation gates, and classes for the qnn, respectively. ‘init’ denotes the initial data distribution for the qnn. ‘data_desc’ denotes the data description. ‘feedback’ denotes the gradient feedback from the previous iteration.
Model architecture of the quantum circuit.
In this study, we evaluate our framework using a backbone qnn consisting of a quantum circuit followed by a fully connected layer. Classical data are first encoded into quantum states via angle encoding, where each feature is mapped to rotation gates (e.g., $R_{X}$) on a specific qubit. This encoding maps data into the Hilbert space while preserving differentiability. The circuit applies repeated layers of parameterized rotations ($R_{X}$, $R_{Y}$, $R_{Z}$) and linear-topology CNOT gates for entanglement. After computation, the quantum state is measured in the computational basis, and expectation values of Pauli-Z operators are computed and used as circuit outputs. These values are then processed by the classical fully connected layer. The overall architecture is adaptable in terms of layers, qubits, and rotation gates, as illustrated in Fig. 13.
Hardware and software.
The experiment is conducted on a server with the following settings:
  • Operating System: Ubuntu 22.04.3 LTS
  • CPU: Intel Xeon w5-3433 @ 4.20 GHz
  • GPU: NVIDIA RTX A6000 48GB
  • Software: Python 3.11.8, PyTorch 2.2.2, Pennylane 0.35.1.
Based on the above computational infrastructure and setup, for example, our search framework can be reproduced in about 15 hours using 18 qubits.
Use of llms.
llms were used only to assist in polishing the language and improving readability. No part of the technical content, analysis, or experimental results was generated by llms.